Gail and Bill drove to a beach at an average speed of 50mph and returned home over the same road at an average speed of 55mph. The trip home took 30 minutes. How far is the beach from their home?
I know that the distance is 27.5 because I know they traveled 55mph in 30minutes. 55/30 = 27.5. I figured this out by the following:
55/60 = d/30 50/60 = 27.5/t
60d = 1650 50t = 1650
d=27.5 miles t = 33 min
but what they are wanting me to do is put the info in a table that represents
rate times time = distance (r)(t) = (d)
such as
r t d
to the beach 50mph
from beach 55mph 30min
Can you help?

Answer by stanbon(48554)

Gail and Bill drove to a beach at an average speed of 50mph and returned home over the same road at an average speed of 55mph. The trip home took 30 minutes. How far is the beach from their home?
=----------------------------
From beach DATA:
rate = 55 mph ; time = (1/2) hr; distance = rt = (55/2) miles = 27.5 miles
------------------------------
To beach DATA:
rate = 50 mph ; time = x hrs; distance = 50x miles
===========================
I don't know if they want it but:
50x = 55/2
x = 0.55 hours (time to get to the beach)
================
Cheers,
Stan H.